# Graph neural networks with configuration cross-attention for tensor compilers

### v1

@@ Abstract
https://github.com/thanhhau097/google_fast_or_slow.
@@ 1.1 Related work
Several attempts have been made to build a highly efficient tensor compiler in recent years. Tensorflow [1] has a rule-based tensor program optimization engine XLA [18] that was studied by [19]. TVM [5] introduces Python-based meta-language to describe the computation and its execution schedule separately, allowing a range of automated optimizations mostly limited to one operator and avoiding operator (kernel) fusion. AutoTVM [6] introduces optimization of tensor programs based on gradient-boosted trees and TreeGRU and uses the ranking loss for model training rather than element-wise losses like MSE. PyTorch [14], being a framework built with the imperative paradigm in mind, in its recent version, supports TorchScript, a just-in-time (JIT) compiled for the annotated functions and classes. JAX [3] as a functional meta-language natively supports JIT.
@@
TASO [9] performs equivalent graph substitution as a way to fuse kernels. PET [21] then builds on top of TASO [9] to expand the search space to non-equivalent transformations and apply automatically generated correction kernels. DeepCuts [11], Ansor [24], and TensorComp [20] rely on heuristics to solve the problem of efficient execution of a computational graph. NN-Meter [23] presents a latency prediction model based on a combination of heuristics to account for the effects of kernel fusion and a random forest for single-operator latency prediction.
@@
All the aforementioned works mostly rely on heuristics and rules to compile a tensor program. While the compilation time of a heuristics-based algorithm may be very small, it fails to achieve the absolute minimum of program runtime. In this work, we propose an algorithm based on machine learning to optimize a tensor program that is represented as a computational graph. The closest work to ours are Phothilimthana, 2020 [15] and Xu, 2023 [22] that use the same dataset and a benchmark TpuGraphs [16]. Graph Segment Training (GST) [4] uses TpuGraphs as well but reports another metric, OPA, and does not provide a breakdown across the collections.
@@ 1.3 Societal impact
We perform a case study to highlight the importance of data center AI workload optimization. According to our estimates the potential impact of this work can be reduction of CO2 emissions equivalent to 50% (or higher) of household emissions in areas similar to North Virginia, VA. The details can be found in Section A.1.
1.4 Dataset and benchmark details
The only publicly available dataset for the large-scale compiler configuration search is TpuGraphs [16]. TpuGraphs contains execution times of an XLA’s HLO graph with a specific compiler configuration on a Tensor Processing Unit (TPU v3). TpuGraphs focuses on optimizing tensor layouts and tensor tiling as compiler configurations. Tensor layout optimization dataset comprises 4 collections organized in a matrix shown in Table 1. The two groups of network architectures (xla and nlp) represent two distinct categories of workloads: xla - predominantly computer vision loads, while nlp - exclusively transformer-based natural language processing loads. Each architecture has up to 100’000 different tensor layout configurations and the associated runtimes recorded. Another dimension across which the layout dataset is organized is the utilized configuration search strategy: random or genetic-algorithm-based (GA-based, denoted as Default). Even though the final goal is to be able to predict configurations’ runtimes, during the dataset creation, some sort of bootstrapping search must be used. Random search gives very wide coverage across all the possible runtimes, whereas the GA-based search focuses more on sampling runtimes in the vicinity of the fastest runtime, making the task of runtime prediction harder and very challenging for the predictive model.
Table 1: The matrix of the 4 Layout collections
Configuration sampling strategy
Random (uniform) Default (GA-based)
NLP (Transformers) layout-nlp-random layout-nlp-default
@@ 2.1 Problem specification
We are looking to find the configuration $\tilde{c}$ that minimizes the tensor program runtime $R(c)$ across the configuration space $C$.
@@ 2.2.1 Graph pruning
For layout collections, only Convolution, Dot, and Reshape nodes are configurable. Also, in most cases, the majority of nodes are identical across the configuration set. Thus, we adopt the following pruning strategy: for each graph, we only keep the nodes that are either configurable nodes themselves or are connected to a configurable node, i.e., input or output to a configurable node. By doing this, we transform a single graph into multiple (possibly disconnected) sub-graphs. The possibly disconnected graph does not pose a problem since TGraph has a global graph pooling layer as one of the final layers that fuses the sub-graph information. This way of graph pruning reduces the vRAM usage 4 times and speeds up training by a factor of 5 in some cases. An example of graph pruning is shown on Figure 1
@@ 2.3 Architecture details
Following the reasoning laid out by [15], we employ GraphSAGE [7] as a basis of a graph convolutional block. GraphSage operation can be expressed as
@@ 2.3.1 Channel-wise self-attention
Inspired by the idea of Squeeze-and-Excitation [8], we add a channel-wise self-attention layer as a part of the graph convolutional block. We first apply a Linear layer to bottleneck the channel dimensions (8x reduction), followed by ReLU. Then, we apply a second linear layer to increase the channels again to the original value, followed by sigmoid. We finish by applying element-wise multiplication to the obtained feature map and the original input. The idea behind channel-wise self-attention is to capture the correlations between channels and use them to suppress less useful ones while enhancing the important ones.
@@ 2.4.1 Loss function
We use the Pairwise Hinge Loss (PairwiseHingeLoss, [10], [2]) loss function for training the model.
@@ 2.4.2 Training details
We use Adam [12] optimizer (specifically AdamW version) with the learning rate of 1e-3, 0.05 of the total number of epochs as linear warm-up, a single-cycle (lifted cosine) learning rate schedule, and weight decay of 1e-5 for non-bias parameters. We apply gradient norm clipping at value 1.0.
@@
Training wall-clock time is 2.5 hours per fold per collection measured on RTX4090 with 24 GB RAM. Training one set of models for all collections produces 13.45 kg CO2 as per [13].
@@ 2.5 Benchmark results
2.5.1 Evaluation splits
TpuGraphs [16] dataset does not provide public test data annotations. Hence, we report the cross-validation score according to the Section 2.4.3.
2.5.2 Evaluation metrics
For the tile collection, the metric is set as:
$M_{tile}=1-\left(\frac{\text{The best runtime of the top-$k$ predictions}}{\text{The best runtime of all configurations}}-1\right)=2-\frac{\min_{i\in K}r_{i}}{\min_{i\in A}r_{i}}$ (10)
where $K=5$.
2.5.3 Details of the inference mode
2.5.4 Experimental results
Table 2: Experimental results
@@ 2.5.4 Experimental results
Collection Metric TpuGraphs [16] Xu 2023 [22] TGraph (ours)
@@
Our experimental results are summarized in the Table 2. The confidence ranges are reported as 1-sigma. We demonstrate state-of-the-art performance in 4 out of 5 collections. On xla-default Xu 2021 [22] show better results than our work; however, their results may contain an error since xla-default collection is harder than xla-random due to closer and harder-to-distinguish runtime annotations (the pattern is also followed by the results of TpuGraphs [16]), but the score of [22] for xla-default is higher than for xla-random which is very implausible.
@@ 3 Conclusion
In general, more efficient ML-based tensor compilation methods have a very positive societal impact. Firstly, they decrease energy consumption and CO2 emissions of data centers, consequently helping to fight climate change. Secondly, they help to free software engineers from the tedious labor of re-implementing lots of highly specialized computational kernels for the constant flow of hardware releases. Even though it may seem that it is a case of "AI taking over people’s jobs", in fact, the achieved extreme efficiency of digital infrastructure like data centers may cover the needs of people to the extent that they do not need to work or can opt to dedicate themselves to more human-centered activities.
@@ 4 Limitations
Acknowledgments and Disclosure of Funding
This work was supported by the SDAIA-KAUST Center of Excellence in Data Science and Artificial Intelligence (SDAIA-KAUST AI).
This work was supported by the prizes of the 1st and the 2nd winning places of "Google - Fast or Slow? Predict AI Model Runtime" Kaggle competition.
Appendix A Appendix / supplemental material
A.1 Environmental impact case study
According to {1} the total data center AI workload consumption in Northern Virginia, VA (NV), the US was 2132 MW in 2023. Thus, the annual data center energy consumption can be estimated as 18.6 million MWh. Considering the carbon footprint of energy production in NV of 0.3 tonne CO2 per MWh as per {2} the total annual CO2 emissions of NV data centers can be assessed as 5.58 mln tonnes CO2. From the authors of XTAT [17] we take 5% as a reference number for the runtime speed-up across a diverse dataset of 150 neural architectures. Speeding up AI workloads by 5% with the more efficient execution would reduce CO2 emissions by 275’000 tonnes CO2 yearly in NV alone. This is equivalent to the annual emissions of 36’000 households (approximately 50% of all NV households). Even though it is yet to be determined how to estimate the real acceleration of computation based on the values of Kendall’s $\tau$, we expect the effect to be similar or superior to XTAT [17].
{1} Angus Loten. Rising Data Center Costs Linked to AI Demands, 2023. (www.wsj.com)
{2} Power sector carbon intensity in the United States in 2022, by state, 2022. (www.statista.com)

### v2

@@ Abstract
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1.1 Related work
Several attempts have been made to build a highly efficient tensor compiler in recent years. Tensorflow Abadi et al. (2016) has a rule-based tensor program optimization engine XLA Sabne (2020) that was studied by Snider & Liang (2023). TVM Chen et al. (2018a) introduces Python-based meta-language to describe the computation and its execution schedule separately, allowing a range of automated optimizations mostly limited to one operator and avoiding operator (kernel) fusion. AutoTVM Chen et al. (2018b) introduces optimization of tensor programs based on gradient-boosted trees and TreeGRU and uses the ranking loss for model training rather than element-wise losses like MSE. PyTorch Paszke et al. (2019), being a framework built with the imperative paradigm in mind, in its recent version, supports TorchScript, a just-in-time (JIT) compiled for the annotated functions and classes. JAX Bradbury et al. (2018) as a functional meta-language natively supports JIT.
@@
TASO Jia et al. (2019) performs equivalent graph substitution as a way to fuse kernels. PET Wang et al. (2021) then builds on top of TASO Jia et al. (2019) to expand the search space to non-equivalent transformations and apply automatically generated correction kernels. DeepCuts Jung et al. (2021), Ansor Zheng et al. (2020), and TensorComp Vasilache et al. (2018) rely on heuristics to solve the problem of efficient execution of a computational graph. NN-Meter Zhang et al. (2021) presents a latency prediction model based on a combination of heuristics to account for the effects of kernel fusion and a random forest for single-operator latency prediction.
@@
All the aforementioned works mostly rely on heuristics and rules to compile a tensor program. While the compilation time of a heuristics-based algorithm may be very small, it fails to achieve the absolute minimum of program runtime. In this work, we propose an algorithm based on machine learning to optimize a tensor program that is represented as a computational graph. The closest work to ours are Phothilimthana et al. (2020) and Xu et al. (2023) that use the same dataset and a benchmark TpuGraphs Phothilimthana et al. (2023). Graph Segment Training (GST) Cao et al. (2023) uses TpuGraphs as well but reports another metric, OPA, and does not provide a breakdown across the collections.
Apart from TpuGraphs, there are few datasets that represent runtime measurements of computational graphs: Tenset Zheng et al. (2021) and the dataset published by the authors of nn-meter Zhang et al. (2021), while none of these explicitly organizes node and edge attributes in a systematic way suitable for machine learning.
1.2 TpuGraphs dataset and benchmark details
Figure 1: An example of how different tensor layout configurations affect the runtime of the computational (sub-)graph. Configuration 1 is faster than and, consequently, superior to configuration 2.
Table 1: The matrix of the 4 Layout collections.
Configuration sampling strategy
Random (uniform) Default (GA-based)
NLP (Transformers) layout-nlp-random layout-nlp-default
The only publicly available dataset for the large-scale compiler configuration search is TpuGraphs Phothilimthana et al. (2023). TpuGraphs contains execution times of an XLA’s HLO graph with a specific compiler configuration on a Tensor Processing Unit (TPU v3). TpuGraphs focuses on optimizing tensor layouts and tensor tiling as compiler configurations. Tensor layout optimization dataset comprises 4 collections organized in a matrix shown in Table 1. The two groups of network architectures (xla and nlp) represent two distinct categories of workloads: xla - predominantly computer vision loads, while nlp - exclusively transformer-based natural language processing loads. Each architecture has up to 100’000 different tensor layout configurations and the associated runtimes recorded. The total number of unique architectures in layout:xla collections is 78 with the average number of configurations of over 11,000 (for layout:xla::random), and in layout:nlp collections - 244 with the average number of configurations of over 66,000 (for layout:nlp::random). Another dimension across which the layout dataset is organized is the utilized configuration search strategy: random or genetic-algorithm-based (GA-based, denoted as Default). Even though the final goal is to be able to predict configurations’ runtimes, during the dataset creation, some sort of bootstrapping search must be used. Random search gives very wide coverage across all the possible runtimes, whereas the GA-based search focuses more on sampling runtimes in the vicinity of the fastest runtime, making the task of runtime prediction harder and very challenging for the predictive model.
To illustrate the problem of configuration selection we provide an example on Figure 1. Here, 4 elementary operations compose a computational graph, while only two of them, reshape and conv, are configurable. A tensor layout can be chosen by the compiler, and the choice results in potentially significantly different runtimes as a result of random or sequential memory access and deep specifics of a particular computational unit. More details can be found in Phothilimthana et al. (2023) Figure 3.
@@ 1.4 Societal impact
We perform a case study to highlight the importance of data center AI workload optimization. According to our estimates the potential impact of this work can be reduction of CO2 emissions equivalent to 50% (or higher) of household emissions in areas similar to North Virginia, VA. The details can be found in Section 2.6.
@@ 2.1 Problem specification
We are looking to find the configuration $\tilde{c}$ that minimizes the tensor program runtime $R(c)$ across the configuration space $C$ for a specific computational graph.
@@
The configuration space $C$ can be described as $\mathbb{Z}^{N}$ where $N$ is the number of discrete configurable variables (node and edge attributes) in a specific graph.
@@ 2.2.1 Graph pruning
For layout collections, only Convolution, Dot, and Reshape nodes are configurable. Also, in most cases, the majority of nodes are identical across the configuration set. Thus, we adopt the following pruning strategy: for each graph, we only keep the nodes that are either configurable nodes themselves or are connected to a configurable node, i.e., input or output to a configurable node. By doing this, we transform a single graph into multiple (possibly disconnected) sub-graphs. The possibly disconnected graph does not pose a problem since TGraph has a global graph pooling layer as one of the final layers that fuses the sub-graph information. This way of graph pruning reduces the vRAM usage 4 times and speeds up training by a factor of 5 in some cases. An example of graph pruning is shown on Figure 2.
@@ 2.3 Architecture details
Following the reasoning laid out by Phothilimthana et al. (2020), we employ GraphSAGE Hamilton et al. (2017) as a basis of a graph convolutional block. GraphSage operation can be expressed as
@@ 2.3.1 Channel-wise self-attention
Inspired by the idea of Squeeze-and-Excitation Hu et al. (2018), we add a channel-wise self-attention layer as a part of the graph convolutional block. We first apply a Linear layer to bottleneck the channel dimensions (8x reduction), followed by ReLU. Then, we apply a second linear layer to increase the channels again to the original value, followed by sigmoid. We finish by applying element-wise multiplication to the obtained feature map and the original input. The idea behind channel-wise self-attention is to capture the correlations between channels and use them to suppress less useful ones while enhancing the important ones.
@@ 2.4.1 Loss function
We use the Pairwise Hinge Loss (PairwiseHingeLoss, Joachims (2002), Agarwal et al. (2019)) loss function for training the model.
@@ 2.4.2 Training details
We use Adam Kingma & Ba (2014) optimizer (specifically AdamW version) with the learning rate of 1e-3, 0.05 of the total number of epochs as linear warm-up, a single-cycle (lifted cosine) learning rate schedule, and weight decay of 1e-5 for non-bias parameters. We apply gradient norm clipping at value 1.0.
@@
Training wall-clock time is 2.5 hours per fold per collection measured on RTX4090 with 24 GB RAM. Training one set of models for all collections produces 13.45 kg CO2 as per Lacoste et al. (2019).
@@ 2.5 Benchmark results
Table 2: Experimental results.
@@
Collection Metric TpuGraphs Phothilimthana et al. (2023) Xu et al. (2023) TGraph (ours)
@@
Table 3: Ablation Study.
Configuration Validation score, Kendall’s $\tau$
layout:xla:random layout:xla:default layout:nlp:random layout:nlp:default
Final, all features 0.6840 0.4785 0.9713 0.5628
- Channel-wise self-attention 0.6737 (-0.0103) 0.4787 (+0.0002) 0.9680 (-0.0033) 0.5555 (-0.0073)
- Cross-configuration attention 0.6539 (-0.0301) 0.4518 (-0.0267) 0.9387 (-0.0326) 0.5436 (-0.0192)
- Graph edges 0.5022 (-0.1818) 0.3631 (-0.1154) 0.7751 (-0.1962) 0.3349 (-0.2279)
2.5.1 Evaluation splits
TpuGraphs Phothilimthana et al. (2023) dataset does not provide public test data annotations. Hence, we report the cross-validation score according to the Section 2.4.3.
2.5.2 Evaluation metrics
For the tile collection, the metric is set as:
$M_{\text{tile}}$ $=1-\left(\frac{\text{Best runtime of top-}k\ \text{predictions}}{\text{Best runtime of all configurations}}-1\right)$
$=2-\frac{\min_{i\in K}r_{i}}{\min_{i\in A}r_{i}}$ (10)
where $K=5$.
2.5.3 Details of the inference mode
2.5.4 Experimental results
Our experimental results are summarized in the Table 2. The confidence ranges are reported as 1-sigma. We demonstrate state-of-the-art performance in 4 out of 5 collections. On xla-default Xu 2021 Xu et al. (2023) show better results than our work; however, their results may contain an error since xla-default collection is harder than xla-random due to closer and harder-to-distinguish runtime annotations (the pattern is also followed by the results of TpuGraphs Phothilimthana et al. (2023)), but the score of Xu et al. (2023) for xla-default is higher than for xla-random which is very implausible.
2.5.5 Ablation study
Ablations for channel-wise self-attention, cross-configuration attention, and edges in the graph are collected in Table 3. While the effect of channel-wise self-attention is less obvious but nevertheless noticeable, the effect of cross-configuration attention is substantial, implying that the task of comparing the configurations between each other is easier than predicting the absolute values of runtimes. Additionally, we ablate the edges of the GraphSage GNN to demonstrate how essential the connectivity between the computational nodes is. In tensor compilers the adjacent operators are often fused into a single optimized operator, the procedure commonly know as kernel fusion. For a model solving the problem of predicting computational graph runtimes it is paramount to implicitly learn the ”rules” of kernel fusion from data since the early stages of tensor compilation including kernel fusion are treated as a black box.
2.6 Environmental impact case study
[!ht] “Data centers will use 8% of US power by 2030, compared with 3% in 2022.” – Goldman Sachs, 2024
According to Loten (2023) the total data center AI workload consumption in Northern Virginia (NV), VA, the US was 2132 MW in 2023. Thus, the annual data center energy consumption can be estimated as 18.6 million MWh. Considering the carbon footprint of energy production in NV of 0.3 tonne CO2 per MWh as per Statista (2022) the total annual CO2 emissions of NV data centers can be assessed as 5.58 mln tonnes CO2. From the authors of XTAT Phothilimthana et al. (2021) we take 5% as a reference number for the runtime speed-up across a diverse dataset of 150 neural architectures. Speeding up AI workloads by 5% with the more efficient execution would reduce CO2 emissions by 275’000 tonnes CO2 yearly in NV alone. This is equivalent to the annual emissions of 36’000 households (approximately 50% of all NV households). Even though it is yet to be determined how to estimate the real acceleration of computation based on the values of Kendall’s $\tau$, we expect the effect to be similar or superior to XTAT Phothilimthana et al. (2021).
@@ 3 Conclusion
In general, more efficient ML-based tensor compilation methods have a very positive societal impact. Firstly, they decrease energy consumption and CO2 emissions of data centers, consequently helping to fight climate change. Secondly, they help to free software engineers from the tedious labor of re-implementing lots of highly specialized computational kernels for the constant flow of hardware releases. Even though it may seem that it is a case of ”AI taking over people’s jobs”, in fact, the achieved extreme efficiency of digital infrastructure like data centers may cover the needs of people to the extent that they do not need to work or can opt to dedicate themselves to more human-centered activities.